OpenStudy (anonymous):

need help with -6+5(-1-b)=19

OpenStudy (anonymous):

If you are trying to solve for b then you just have to get b then you need to get b by itself. -6+5(-1-b)=19 First you have to distribute the 5 to the (-1-b) -6-5-5b=19 Then you need to add all your like terms -11-5b=19 Now you need to get b by itself by adding 11 to both sides. -5b=30 Now if you are multiplying now, you must do the oppotiste. Divide to get b by itself. That will give you your answer.